Jeremana72:

Solve for x: (5 points) negative 3 over 2, multiplied by x minus 9 equals negative 27 −24 −12 11 12

HaydenwH:

have you wrote it out??

Jeremana72:

yes

HaydenwH:

have you attempted to solve it??

Jeremana72:

yes

HaydenwH:

what did you come up with

Jeremana72:

guide me thru how to do it plz

Jeremana72:

plz

HaydenwH:

ok one sec

Jeremana72:

someone actually started helping me on brainly bbye

Hoodmemes:

Ok bro so first ya wanna -3x/2 - 9 = -27, -3x/2 = -27 + 9, -3x/2 = -18, -3x = -18 * 2, -3x = -36, x = 36/3, x = (blank)

HaydenwH:

dmn my bad for tryin work it out